### v1

# ViP2-CLIP: Visual-Perception Prompting with Unified Alignment for Zero-Shot Anomaly Detection

Abstract
Zero-shot anomaly detection (ZSAD) aims to detect anomalies without any target domain training samples, relying solely on external auxiliary data. Existing CLIP-based methods attempt to activate the model’s ZSAD potential via handcrafted or static learnable prompts. The former incur high engineering costs and limited semantic coverage, whereas the latter apply identical descriptions across diverse anomaly types, thus fail to adapt to complex variations. Furthermore, since CLIP is originally pretrained on large-scale classification tasks, its anomaly segmentation quality is highly sensitive to the exact wording of class names, severely constraining prompting strategies that depend on class labels. To address these challenges, we introduce ViP2-CLIP. The key insight of ViP2-CLIP is a Visual-Perception Prompting (ViP-Prompt) mechanism, which fuses global and multi-scale local visual context to adaptively generate fine-grained textual prompts, eliminating manual templates and class-name priors. This design enables our model to focus on precise abnormal regions, making it particularly valuable when category labels are ambiguous or privacy-constrained. Extensive experiments on 15 industrial and medical benchmarks demonstrate that ViP2-CLIP achieves state-of-the-art performance and robust cross-domain generalization.
ViP2-CLIP: Visual-Perception Prompting with Unified Alignment for Zero-Shot Anomaly Detection
Ziteng Yang1,*    JingzehuaXu1,*    Yanshu Li2    Zepeng Li1 Yeqiang Wang3    Xinghui Li1$\dagger$ 1Tsinghua Shenzhen International Graduate School 2Brown University    3Northwest A&F University
Figure 1: Comparison between prior CLIP-based methods and ViP2-CLIP. ViP2-CLIP introduces ViP-Prompt to replace fixed class-name tokens with image-conditioned prompts that fuse global and local cues, and employs UTPA to perform a unified patch-level alignment.
1 Introduction
Large-scale vision–language models (VLMs) such as CLIP (Radford et al., 2021) have shown impressive zero-shot recognition capabilities by pre-training on vast image–text pairs. Inspired by this success, CLIP has been widely adopted for Zero-Shot Anomaly Detection (ZSAD) (Jeong et al., 2023), which seeks to detect anomalies automatically without any target domain training samples, offering a practical solution for data-scarce scenarios. Unlike conventional image classification, which focuses on foreground semantics, anomaly detection emphasizes irregular regions on object surfaces. Consequently, existing CLIP-based ZSAD methods typically construct two sets of text prompts—‘normal’ and ‘anomalous’, and align them with image features in a shared semantic space. As illustrated in Figure 1(a), text prompts are aligned with the global image feature to infer image-level anomaly score, while alignment with local patch features then generates pixel-level anomaly maps.
Currently, state-of-the-art ZSAD methods focus on prompt engineering to leverage CLIP’s generalization. Representative approaches such as WinCLIP (Jeong et al., 2023), AnomalyCLIP (Zhou et al., 2024), and VCP-CLIP (Qu et al., 2024) either craft semantically rich manual templates or learn prompt tokens to boost performance. However, manual templates demand extensive human effort and cannot guarantee comprehensive semantic coverage. Although some works (Zhou et al., 2024) introduce object-agnostic learnable prompts to capture generic normal and abnormal concepts, they apply the same description across diverse anomaly types, hindering adaptability to complex semantic variations. Moreover, methods like CLIP-AD (Chen et al., 2024), AdaCLIP (Cao et al., 2024), FILo (Gu et al., 2024), and KANOCLIP (Li et al., 2025) rely heavily on class names as prompt priors. Yet we find that CLIP’s segmentation quality remains acutely sensitive to class-name variations, even after fine-tuning on auxiliary anomaly data (see Appendix A for detailed analysis). This likely stems from the limited scope of downstream adaptation, which cannot fully stabilize local alignment against label shifts. Such sensitivity constrains the class-name usage in prompts, highlighting the urgent need for robust, class-name agnostic but adaptive prompting strategies in ZSAD.
To this end, we propose a Visual-Perception Prompting (ViP-Prompt) mechanism composed of an Image-Conditioned Adapter (ICA) and a Fine-Grained Perception Module (FGP). ICA adaptively injects global visual context into the prompts’ embedding space, while FGP integrates multi-scale patch features, enabling the prompts to capture fine-grained irregularities. By replacing fixed class-name tokens with image-conditioned prompts that fuse global and local cues, ViP-Prompt significantly enhances prompts’ robustness and reinforces cross-modal alignment across diverse semantic scenarios. Building upon that, we further introduce ViP2-CLIP, a ZSAD model that integrates visual-perception prompts with a unified alignment mechanism; its overall architecture is depicted in Figure 1(b). ViP2-CLIP first employs ViP-Prompt to condition text prompts on both global and local visual cues, fully leveraging CLIP’s ZSAD performance; it then adopts a Unified Text-Patch Alignment (UTPA) strategy to align prompt tokens with multi-scale patch features through a single optimization, yielding superior precision in both detection and localization.
Overall, our key contributions are summarized as follows:
  • We propose ViP-Prompt, which fuses global and local visual features to generate image-conditioned normal and abnormal prompts, fully exploiting CLIP’s potential for ZSAD. This design eliminates the need for handcrafted templates or class-name priors, delivers stronger generalization and robustness, making it particularly valuable when category labels are ambiguous or privacy-constrained.
  • On top of that, we present ViP2-CLIP, which integrates a simple but effective strategy-UTPA, that for the first time, introduces a consistency-alignment strategy within training-based CLIP models, enhancing CLIP’s performance in both anomaly detection and localization.
  • Comprehensive experiments on 15 real-world industrial and medical datasets demonstrate that ViP2-CLIP consistently achieves superior ZSAD performance.
2 Related Work
Zero-Shot Anomaly Detection (ZSAD)
Zero-shot anomaly detection (ZSAD) localizes abnormal regions without any target class samples during training. The advent of CLIP (Radford et al., 2021) has spurred notable advances in this task. WinCLIP (Jeong et al., 2023) first ported CLIP to ZSAD with diverse handcrafted prompts. However, since CLIP is pretrained to align with general object semantics rather than abnormal patterns, its anomaly detection capability remains limited. To ease this burden, APRIL-GAN (Chen et al., 2023) and CLIP-AD (Chen et al., 2024) append shallow linear heads and fine-tune them on auxiliary datasets. AnomalyCLIP (Zhou et al., 2024) removes handcrafted templates via learnable, object-agnostic prompts, boosting cross-domain generalization, while AdaCLIP (Cao et al., 2024) employs hybrid prompts to align text and image features better. VCP-CLIP (Qu et al., 2024) instead projects the image’s global feature into a fixed handcrafted prompt space, which enhances segmentation performance. By contrast, our ViP-Prompt fuses multi-scale visual cues with a learnable seed to generate dynamic, manual-free prompts, fully exploiting CLIP’s ZSAD performance, boosting both effectiveness and robustness across diverse domains.
Prompt Learning
Prompt learning was initially proposed in natural language processing to enhance the adaptability of pre-trained models (Ouyang et al., 2022; Touvron et al., 2023) to diverse downstream tasks. CoOp (Radford et al., 2021) first brought this paradigm to the vision domain by inserting trainable tokens into text inputs, allowing CLIP to adapt to specific tasks without fine-tuning. Static prompts, however, soon proved brittle on unseen classes. To enhance transferability, CoCoOp (Zhou et al., 2022) and DenseCLIP (Rao et al., 2022) generate image-conditioned prompts that respond to visual context. Later works such as MaPLe (Khattak et al., 2023a) and PromptSRC (Khattak et al., 2023b) push prompting into both modalities, jointly tuning image and text space for more substantial cross-modal alignment. More recent works (Yao et al., 2023, 2024) enforce prompt generalization by regularizing learnable tokens toward handcrafted templates, reducing forgetting and improving unseen class recognition.
Figure 2: Framework of ViP2-CLIP. ViP2-CLIP first introduces ViP-Prompt to enhance cross-modal alignment: ViP-ICA injects global visual context into the prompts’ embedding space, while ViP-FGP fuses local patch features to enhance the prompts’ fine-grained perceptual capacity. Finally, the UTPA module performs unified alignment in multiple layers to jointly support image-level anomaly detection and pixel-level anomaly localization.
3 Preliminary
CLIP consists of a text encoder $T(\cdot)$ and a visual encoder $F(\cdot)$, both implemented as mainstream multi-layer networks. Leveraging contrastive learning on large-scale image–text pairs, CLIP achieves superior zero-shot recognition. Given a class name $c$, we combine it with a text prompt template $G$ (e.g., ‘A photo of a [cls]’, where [cls] represents the category class-name). The resulting text is then fed into the text encoder $T(\cdot)$ to obtain the prompt embedding $g_{c}=T\bigl{(}G(c)\bigr{)}\in\mathbb{R}^{D}$. For an input image $x_{i}$, the visual encoder generates its global visual embedding $f_{i}\in\mathbb{R}^{D}$ and the local patch embeddings $f_{i}^{m}\in\mathbb{R}^{H\times W\times D}$. Specifically, given a category set $C$, CLIP computes the probability of image $x_{i}$ belonging to $c$ as follows:
${}{p\bigl{(}y=c\mid{x_{i}}\bigr{)}=P(g_{c},f_{i})=\frac{\exp\!\bigl{(}\langle g_{c},f_{i}\rangle/\tau\bigr{)}}{\sum_{c^{\prime}\in C}\exp\!\bigl{(}\langle g_{c^{\prime}},f_{i}\rangle/\tau\bigr{)}}},$ (1)
[…]
Unlike conventional classification, ZSAD flags deviations from normality rather than assigning foreground semantics. Most approaches, therefore, instantiate two text prompts Jeong et al. (2023), a normal prompt $g_{n}$ and an abnormal prompt $g_{a}$. At the image level, they compare the global image feature $f_{i}$ with both prompts and take $P(g_{a},f_{i})$ as the anomaly score; while at the pixel level, for each location $(j,k)$, they extract the patch token $f_{i}^{m}(j,k)$, and compute the local normal score $S_{n}(j,k)=P\bigl{(}g_{n},f_{i}^{m}(j,k)\bigr{)}$ and anomaly score $S_{a}(j,k)=P\bigl{(}g_{a},f_{i}^{m}(j,k)\bigr{)}$, from which a pixel-wise anomaly map is generated.
4 Method
We propose ViP2-CLIP, which leverages Visual-Perception Prompting (ViP-Prompt) to fully exploit CLIP’s ZSAD capabilities. As illustrated in Figure 2, ViP2-CLIP first introduces ViP-Prompt to fuse learnable normal and abnormal prompts with both global and multi-scale local visual features. This design enables the prompts to adaptively capture the fine-grained visual patterns of the inspected image, thereby improving cross-modal semantic alignment. Furthermore, we adopt a Unified Text-Patch Alignment (UTPA) scoring strategy, which optimizes alignment between prompt tokens and multi-scale patch features to support anomaly detection and localization jointly. By aggregating alignment signals across multiple layers, ViP2-CLIP produces robust anomaly scores and fine-grained anomaly maps.
4.1 Visual-Perception Prompt (ViP-Prompt)
Effective prompt design plays a crucial role in achieving strong ZSAD performance. Handcrafted templates (Jeong et al., 2023; Chen et al., 2023) are costly and fragile, while static learnable prompts (Zhou et al., 2024) apply identical textual descriptions to all detection targets, limiting their ability to adapt to complex semantic variations. Furthermore, most existing techniques (Chen et al., 2024; Cao et al., 2024; Gu et al., 2024; Li et al., 2025) generate prompts based on class names, since CLIP’s anomaly segmentation quality is highly sensitive to class-name wording, these methods remain unusable in label-ambiguous scenarios. To overcome these limitations, we therefore introduce ViP-Prompt, which is built from an Image-Conditioned Adapter (ICA) and Fine-Grained Perception Module (FGP). By fusing learnable tokens with the global and local visual context, ViP-Prompt generates multi-level descriptions that flexibly track objects’ visual patterns, thus markedly exploiting CLIP’s ZSAD performance in diverse scenarios.
Image-Conditioned Adapter (ICA)
To eliminate dependence on handcrafted templates and class-name priors, we first define static learnable prompt templates for both normal and anomaly classes:
$g_{n}=[v_{1}][v_{2}]\dots[v_{e}]\;\text{good}\;[o_{1}]\dots[o_{r}],$ (2)
$g_{a}=[w_{1}][w_{2}]\dots[w_{e}]\;\text{damaged}\;[o_{1}]\dots[o_{r}],$ (3)
here, $\{v_{i}\},\{w_{i}\}\in\mathbb{R}^{C}$ denote the learnable normal and anomaly vector, respectively. $\{o_{i}\}\in\mathbb{R}^{C}$ is a generic learnable token that replaces the explicit class label. We adopt the adjectives ‘good’ and ‘damaged’ to guide the prompts to learn richer normal and anomalous semantics. By autonomously learning $g_{n}$ and $g_{a}$, the model can capture generic anomaly patterns across diverse objects.
To condition prompts on the target object and thus adapt to complex detect scenarios, we map the global visual embedding $f$ extracted by the image encoder through a lightweight Meta-Net $h_{i}(\cdot)$ (Linear–ReLU–Linear) into the text embedding space, denoted as $\{a_{i}\}_{i=1}^{r}=h_{i}(f)$, where $r$ is the number of mapped tokens and $a_{i}\in\mathbb{R}^{C}$, then fuse it with static learnable vector to generate dynamic prompt token. Thus, we define the prompts’ structure as follows:
$g_{n}=[v_{1}][v_{2}]\dots[v_{e}]\;\text{good}\;[z_{1}]\dots[z_{r}],$ (4)
$g_{a}=[w_{1}][w_{2}]\dots[w_{e}]\;\text{damaged}\;[z_{1}]\dots[z_{r}],$ (5)
where $z_{i}=o_{i}+a_{i}$. By projecting the image’s global feature into text embedding space, ICA adaptively generates normal and anomalous descriptors conditioned on the target object, eliminating the dependence on explicit class labels and enabling more robust cross-modal alignment in privacy-constrained settings.
Fine-Grained Perception Module (FGP)
To further enhance the prompts’ fine-grained perceptual capacity, we introduce an attention-based interaction module between the text prompts and multi-scale local image features. Specifically, we project both text embeddings and local patch embeddings into a shared $C$-dimensional space. Let $G_{t}\in\mathbb{R}^{2\times C}$ be the prompt embeddings from the text encoder, and $F_{m}^{l}\in\mathbb{R}^{HW\times D}$ be the smoothed local visual embeddings from the $l$-th layer of the visual encoder, we then apply three learnable linear mappings to produce query $Q_{t}=G_{t}W_{q}$, key $K_{m}^{l}=F_{m}^{l}W_{k}$, and value $V_{m}^{l}=F_{m}^{l}W_{v}$, where $W_{q}\in\mathbb{R}^{C\times C}$, $W_{k},W_{v}\in\mathbb{R}^{D\times C}$, are learnable projection matrices.
Then, we compute scaled dot-product attention via $\mathrm{Softmax}\bigl{(}Q_{t}K_{m}^{l\top}/\sqrt{C}\bigr{)}$, and apply these weights to the values $V_{m}^{l}$, yielding enhanced token embeddings. Finally, a linear projection $W_{o}\in\mathbb{R}^{C\times D}$ transforms these embeddings into the layer-specific prompts $\tilde{G_{t}^{l}}$, which integrate fine-grained local features:
$\tilde{G_{t}^{l}}=\bigl{(}\mathrm{Softmax}(Q_{t}K_{m}^{l\top}/\sqrt{C})\,V_{m}^{l}\bigr{)}\,W_{o},$ (6)
here, $\tilde{G_{t}^{l}}\in\mathbb{R}^{2\times D}$ comprises the fine-grained normal prompt $\tilde{g_{n}}$ and anomalous prompt $\tilde{g_{a}}$.
To assess FGP’s fine-grained sensitivity, Figure 3 visualizes the attention maps of both normal and abnormal prompts across diverse images. Normal prompts activate on regular textures and structures of the image, while anomalous prompts concentrate on defects and irregular regions, indicating high sensitivity to anomalous patterns. These results demonstrate that FGP effectively refines prompt embeddings by incorporating localized visual cues, thereby enhancing cross-modal alignment.
4.2 Unified Text-Patch Alignment (UTPA)
Recent CLIP-based ZSAD models typically adopt a dual-branch alignment scheme Chen et al. (2024); Zhou et al. (2024); Chen et al. (2023): one branch aligns the text prompt with global visual embedding to score anomalies at the image level, while the other aligns it with local patch embeddings for pixel-level anomaly maps. However, because these branches operate in separate feature subspaces, it hinders the simultaneous attainment of high-precision anomaly detection and localization.
Figure 3: The visualization result of the attention maps from different prompts in our FGP module.
To address this limitation, we introduce the UTPA strategy, which treats anomaly detection and localization as a unified alignment task between text embeddings and local patch embeddings. Concretely, to obtain the pixel-level anomaly map, we follow the standard practice: for the $l$-th layer feature map $F_{m}^{l}$, the patch token at spatial location $(j,k)$, denoted as $f_{j,k}^{l}\in\mathbb{R}^{D}$, receives its anomaly score by computing the similarity $P\bigl{(}\tilde{g_{a}},\,f_{j,k}^{l}\bigr{)}$ as follows:
${}{P\bigl{(}\tilde{g_{a}},\,f_{j,k}^{l}\bigr{)}=\frac{\exp\!\bigl{(}\langle\tilde{g_{a}},\,f_{j,k}^{l}\rangle/\tau\bigr{)}}{\exp\!\bigl{(}\langle\tilde{g_{n}},\,f_{j,k}^{l}\rangle/\tau\bigr{)}+\exp\!\bigl{(}\langle\tilde{g_{a}},\,f_{j,k}^{l}\rangle/\tau\bigr{)}}}.$ (7)
Next, we upsample the anomaly scores of all patch tokens to produce the anomaly map $M_{a}^{l}$:
$M_{a}^{l}=\mathrm{Up}\bigl{(}P(\tilde{g_{a}},\,F_{m}^{l})\bigr{)},$ (8)
where $Up(\cdot)$ denotes channel-wise upsampling.
Unlike prior works (Chen et al., 2024; Zhou et al., 2024; Chen et al., 2023) that rely on global alignment, we contend that an image’s abnormality is driven by its most defective regions. Consequently, our image-level score is still derived from local patch features. For the $l$-th layer feature map $F_{m}^{l}$, we select the Top-k patch tokens with the highest $P\bigl{(}\tilde{g_{a}},\,f_{j,k}^{l}\bigr{)}$ among all patches. The image’s representative anomaly feature $F_{a}^{l}$ is then obtained via global average pooling over this subset:
$F_{a}^{l}=\frac{1}{k}\sum_{(j,k)\in\mathcal{T}_{l}}f^{l}_{j,k},$ (9)
where $\mathcal{T}_{l}$ indexes the chosen Top-k patches. The Top-k pooling directs the model’s attention to the most suspicious patches, ensuring that the subsequent anomaly score faithfully reflects the severity of localized defects. Consequently, the image-level anomaly score $S_{a}^{l}$ at the $l$-th layer is obtained by computing $S_{a}^{l}=P\bigl{(}\tilde{g_{a}},\,F_{a}^{l}\bigr{)}$.
Figure 4: Visualization of anomaly maps of different ZSAD methods. Our proposed ViP2-CLIP achieves the sharpest segmentations, capturing fine-grained defects in both industrial and medical datasets.
4.3 Training and Inference
During training, ViP2-CLIP is optimized by minimizing a global loss $L_{\mathrm{global}}$ and a local loss $L_{\mathrm{local}}$:
$L_{\mathrm{total}}=L_{\mathrm{global}}+\lambda\sum_{l=1}^{N}L_{\mathrm{local}}^{M^{l}},$ (10)
here, $\lambda$ is a hyperparameter to balance the global and local losses; $N$ represents the number of intermediate layers. $L_{\mathrm{global}}$ is the cross-entropy loss matching the cosine similarity between text embeddings $\tilde{g_{n}}/\tilde{g_{a}}$ and the representative anomalous visual feature $F_{a}^{l}$; $L_{\mathrm{local}}$ is computed as the sum of a Focal loss Lin et al. (2017) and a Dice loss Li et al. (2019) to optimize local alignment jointly:
$\begin{split}L_{\mathrm{local}}=\mathrm{Focal}\bigl{(}\mathrm{Up}([M_{n}^{l},M_{a}^{l}]),S_{\mathrm{gt}}\bigr{)}\\ +\mathrm{Dice}\bigl{(}\mathrm{Up}(M_{n}^{l}),\,I-S_{\mathrm{gt}}\bigr{)}\\ +\mathrm{Dice}\bigl{(}\mathrm{Up}(M_{a}^{l}),\,S_{\mathrm{gt}}\bigr{)},\end{split}$ (11)
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes channel-wise concatenation; for each layer $l$, $M_{n}^{l}$ and $M_{a}^{l}$ are the normal and anomaly score maps; $S_{\mathrm{gt}}$ is the ground-truth mask; and $I$ is the all-ones matrix.
During inference, the final image-level anomaly score is computed as:$\mathrm{Score}=\frac{1}{N}\sum_{l=1}^{N}S_{a}^{l}$. For pixel-level prediction, we merge all intermediate maps $M_{n}^{l}$ and $M_{a}^{l}$, then obtain the anomaly map $\mathrm{Map}\in\mathbb{R}^{H\times W}$ as follows:
${\mathrm{Map}=G_{\sigma}\!\Biggl{(}\frac{1}{N}\sum_{l=1}^{N}\Bigl{(}\tfrac{1}{2}\bigl{(}I-\mathrm{Up}(M_{n}^{l})\bigr{)}+\tfrac{1}{2}\,\mathrm{Up}(M_{a}^{l})\Bigr{)}\Biggr{)}},$ (12)
where $G_{\sigma}$ denotes Gaussian smoothing.
5 Experiments
5.1 Setup
Datasets & Baselines
To comprehensively evaluate ViP2-CLIP across diverse application scenarios, we conduct extensive experiments on public benchmarks spanning industrial and medical domains. In the industrial domain, we use MVTec AD (Bergmann et al., 2019), VisA (Zou et al., 2022), MPDD (Jezek et al., 2021), BTAD (Mishra et al., 2021), KSDD (Tabernik et al., 2020), DAGM (Wieler and Hahn, 2007), and DTD-Synthetic (Aota et al., 2023). For medical imaging, we include brain tumour detection datasets HeadCT (Salehi et al., 2021), BrainMRI (Salehi et al., 2021), and Br35H (Hamada, 2020), skin cancer detection dataset ISIC Codella et al. (2018), colon polyp detection datasets CVC-ClinicDB Bernal et al. (2015), CVC-ColonDB (Tajbakhsh et al., 2015), Endo (Hicks et al., 2021), and Kvasir (Jha et al., 2019). We compare ViP2-CLIP against five leading ZSAD methods: CLIP (Radford et al., 2021), WinCLIP (Jeong et al., 2023), APRIL-GAN Chen et al. (2023), AnomalyCLIP (Zhou et al., 2024) and AdaCLIP (Cao et al., 2024). More details about the datasets or methods are shown in Appendix B.
Task Datasets $|\mathcal{C}|$ CLIP WinCLIP APRIL-GAN AnomalyCLIP AdaCLIP ViP2-CLIP
OpenCLIP CVPR 2023 CVPRw 2023 ICLR 2024 ECCV 2024 -
Image-level (AUROC, AP, F1) MVTec AD 15 (66.5, 82.6, 86.7) (91.8, 96.5, 92.7) (86.1, 93.5, 90.4) (91.6, 96.4, 92.7) (90.1, 95.6, 92.3) (91.2, 96.0, 92.0)
VisA 12 (60.2, 66.2, 73.0) (78.1, 81.2, 78.2) (77.5, 80.9, 78.7) (82.0, 85.3, 80.4) (87.2, 89.7, 83.5) (88.5, 90.4, 84.8)
MPDD 6 (58.7, 69.7, 76.0) (61.5, 69.2, 77.5) (76.8, 83.0, 81.0) (77.5, 82.5, 80.4) (74.8, 78.6, 83.3) (79.7, 84.5, 82.4)
KSDD 1 (74.3, 55.3, 57.8) (92.4, 82.9, 77.7) (96.5, 91.2, 85.4) (97.8, 94.2, 89.7) (97.2, 92.6, 89.5) (98.1, 95.8, 93.2)
BTAD 3 (25.7, 49.8, 66.0) (68.2, 70.9, 67.8) (73.7, 69.7, 68.2) (88.2, 88.2, 83.8) (89.3, 96.5, 90.9) (95.0, 98.4, 94.7)
DAGM 10 (55.3, 43.2, 49.6) (91.8, 79.5, 75.7) (94.4, 83.9, 80.2) (97.7, 92.4, 90.1) (98.2, 92.3, 90.9) (98.5, 94.3, 92.6)
DTD-Synthetic 12 (52.3, 77.8, 85.1) (95.1, 97.7, 94.1) (85.6, 94.0, 89.1) (93.9, 97.2, 93.6) (96.3, 98.1, 95.5) (95.5, 98.1, 94.3)
AVERAGE - (56.1, 63.5, 70.6) (82.1, 82.1, 80.5) (84.4, 85.2, 81.9) (89.8, 90.9, 87.2) (90.4, 91.9, 89.4) (92.4, 93.9,90.6)
Pixel-level (AUROC, PRO, F1) MVTec AD 15 (35.6, 10.6, 6.9) (85.1, 64.6, 24.8) (87.6, 44.0, 43.3) (91.1, 81.4, 39.1) (89.6, 37.8, 45.1) (90.5, 87.1, 43.1)
VisA 12 (43.6, 14.0, 1.5) (79.6, 56.8, 9.0) (94.2, 86.6, 32.3) (95.5, 86.7, 28.3) (95.5, 56.8, 37.0) (95.4, 92.2, 33.6)
MPDD 6 (56.2, 27.3, 8.2) (71.2, 40.5, 15.4) (94.3, 83.8, 31.3) (96.5, 88.7, 34.2) (96.1, 60.3, 31.9) (97.2, 92.6, 35.9)
KSDD 1 (21.7, 2.6, 0.5) (92.8, 70.3, 15.8) (93.2, 84.1, 43.6) (98.1, 94.9, 56.5) (98.4, 53.0, 52.8) (98.5, 96.2, 53.0)
BTAD 3 (41.2, 10.8, 6.8) (72.7, 27.5, 18.5) (89.3, 68.7, 40.6) (94.2, 75.4, 49.7) (90.7, 22.3, 51.4) (95.6, 86.1, 52.7)
DAGM 10 (49.6, 12.4, 2.4) (87.6, 65.7, 12.7) (82.4, 66.0, 37.4) (95.6, 91.0, 58.9) (94.3, 42.5, 59.6) (97.5, 95.2, 61.2)
DTD-Synthetic 12 (43.8, 16.0, 3.7) (79.5, 51.4, 16.1) (95.2, 87.3, 67.4) (97.9, 92.0, 62.2) (98.5, 75.0, 71.8) (99.0, 96.5, 67.5)
AVERAGE - (41.7, 13.4, 4.3) (79.9, 52.6, 16.0) (90.9, 74.4, 42.2) (95.6, 87.2, 47.0) (94.7, 49.7, 49.9) (96.2,92.3, 49.6)
Table 1: ZSAD performance comparison on industrial domain. * denotes results taken from original papers. The best performance is highlighted in red, and the second in blue.
Task Datasets $|\mathcal{C}|$ CLIP WinCLIP APRIL-GAN AnomalyCLIP AdaCLIP ViP2-CLIP
OpenCLIP CVPR 2023 CVPRw 2023 ICLR 2024 ECCV 2024 -
Image-level (AUROC, AP, F1) HeadCT 1 (67.8, 62.4, 70.9) (81.8, 80.2, 78.9) (89.1, 89.4, 82.1) (93.0, 91.1, 88.4) (94.0, 91.4, 90.1) (94.3, 93.9, 88.1)
BrainMRI 1 (72.2, 81.5, 76.5) (86.6, 91.5, 84.1) (89.4, 91.0, 88.2) (90.0, 92.1, 86.5) (94.3, 95.5, 92.2) (95.3, 96.7, 92.3)
Brain35H 1 (76.3, 77.7, 72.2) (79.9, 82.2, 74.0) (91.6, 92.1, 84.5) (93.4, 93.8, 86.4) (95.7, 95.8, 91.1) (95.8, 96.0, 90.1)
AVERAGE - (72.1, 73.9, 73.2) (82.8, 84.6, 79.0) (90.0, 90.8, 84.9) (92.1, 92.3, 87.1) (94.7, 94.2, 91.1) (95.1, 95.5, 90.2)
Pixel-level (AUROC, PRO, F1) ISIC 1 (43.2, 7.6, 44.0) (83.3, 55.1, 64.1) (89.4, 77.2, 71.4) (89.4, 78.4, 71.6) (91.3, 53.2, 75.5) (90.3, 82.3, 73.7)
CVC-ColonDB 1 (59.8, 30.4, 17.8) (64.8, 28.4, 21.0) (78.4, 64.6, 29.7) (81.9, 71.2, 37.5) (81.5, 64.3, 33.6) (82.5, 74.8, 36.2)
CVC-ClinicDB 1 (63.1, 33.8, 23.4) (70.3, 32.5, 27.2) (80.5, 60.7, 38.7) (82.9, 68.1, 42.4) (83.9, 65.7, 42.3) (86.3, 72.1, 44.9)
Endo 1 (59.7, 25.2, 28.9) (68.2, 28.3, 32.9) (81.9, 54.9, 44.8) (84.2, 63.4, 50.3) (86.0, 63.4, 51.8) (84.5, 63.7, 49.2)
Kvasir 1 (58.0, 18.3, 29.3) (69.7, 24.5, 35.9) (75.0, 36.3, 40.0) (79.0, 45.4, 46.2) (81.6, 49.1, 47.1) (81.9, 47.8, 47.3)
AVERAGE - (56.8, 23.1, 28.7) (71.3, 33.8, 36.2) (81.0, 58.7, 44.9) (83.5, 65.3, 49.6) (84.9, 59.1, 50.1) (85.1, 68.1, 50.3)
Table 2: ZSAD performance comparison on medical domain. The best performance is highlighted in red, and the second in blue. Note that the image-level medical AD datasets do not provide pixel-level segmentation annotations, so the pixel-level medical AD datasets are different from the image-level datasets.
Metrics
We adopt standard metrics for evaluation: for image-level anomaly detection, we report the Area Under the Receiver Operating Characteristic curve (AUROC), Average Precision (AP), and maximum F1 score (F1); for pixel-level anomaly segmentation, we report AUROC, the Area Under the Per-Region Overlap curve (AUPRO), and F1.
Implementation Details
We use the publicly released CLIP (ViT-L/14@336px) as the frozen backbone. The learnable prompts contain 10 tokens, 3 of which interact explicitly with the global visual embedding. In UTPA, we retain the top 50 anomalous patches at each alignment layer as image-level descriptors and, following prior work Zhou et al. (2024); Cao et al. (2024), apply cross-modal alignment at layers 6, 12, 18, and 24. We fine-tune ViP2-CLIP on MVTec AD’s test split and evaluate zero-shot performance on all other datasets; for MVTec AD, we fine-tune on VisA’s test set. Dataset-level metrics are averaged across all subclasses. All experiments are implemented in PyTorch 2.6.0 on a single NVIDIA L20 (42 GB). Further implementation details are provided in Appendix B.
5.2 Main Results
Zero-Shot Anomaly Detection on Industrial Datasets
Table 1 compares ViP2-CLIP against five representative baselines on seven industrial defect benchmarks. Our model delivers competitive results and outperforms all baselines in most datasets. The vanilla CLIP struggles, as its pre-training emphasizes general object semantics rather than anomaly patterns. WinCLIP and APRIL-GAN improve detection through handcrafted prompts and local features tuning. AnomalyCLIP employs object-agnostic prompts but remains limited by static prompt design and shallow cross-modal fusion. AdaCLIP further refines prompts in both visual and textual space, which boosts F1 but neglects AUPRO, leading to an imbalanced ZSAD. In contrast, ViP2-CLIP yields consistent improvements across all metrics, achieving state-of-the-art performance. Our ViP-Prompt creates fine-grained normal and abnormal prompts that align tightly with object features; paired with UTPA’s unified alignment, it markedly boosts accuracy at both image-level detection and pixel-level localization.
Figure 4 compares our approach with four strong baselines, illustrating that ViP2-CLIP better captures defects with sharper boundaries, corroborating the quantitative improvements. We also provide computational efficiency analyses in Appendix C.
Zero-Shot Anomaly Detection on Medical Datasets
To assess cross-domain transferability, we benchmark ViP2-CLIP on eight public medical defect datasets. As shown in Table 2, ViP2-CLIP achieves consistently superior detection performance on most medical datasets, generally outperforming prior methods. Qualitative examples in Figure 4 further support this claim: our model precisely localizes melanoma lesions in dermoscopy images and colonic polyps in endoscopic frames, demonstrating robust adaptability to disparate pathological patterns.
Static prompts vs. Visual-Perception prompts
To quantify the benefit of visual conditioning, we compare ViP2-CLIP against ViP2-CLIP_re, which only uses static learnable prompts without visual cues. Figure 5 reports image-level and pixel-level F1 gains on the seven industrial datasets. ViP2-CLIP achieves significantly higher pixel-level F1 on all datasets and improves image-level F1 on five of them, confirming that injecting global and local visual context into prompts markedly strengthens cross-modal alignment, thus boosting superior ZSAD performance.
Figure 5: F1 gains of using visual-conditioned prompts compared to static learnable prompts.
Module MVTec AD VisA
Pixel-level Image-level Pixel-level Image-level
Base (37.8, 11.5, 7.0) (74.1, 87.6, 87.2) (43.5, 14.6, 2.8) (60.2, 66.3, 74.5)
+G1 (89.5, 79.9, 38.6) (71.8, 86.2, 88.5) (95.2, 90.9, 31.6) (79.7, 82.8, 79.5)
+G2 (87.7, 83.8, 42.4) (88.9, 94.9, 91.6) (95.2, 91.6, 35.5) (83.0, 85.9, 81.4)
+G3 (90.5, 87.1, 43.1) (91.2, 96.0, 92.0) (95.4, 92.2, 33.6) (88.5, 90.4, 84.8)
Table 3: Module ablation.
5.3 Ablation Study
Module Ablation
We isolate the impact of three key components: ViP-ICA (G1), ViP-FGP (G2), and UTPA (G3). Table 3 shows that adding each module yields a clear, consistent gain, validating their standalone effectiveness. Specifically, ICA adaptively injects global visual context into prompt embeddings, while FGP incorporates fine-grained local cues, enabling the prompts to capture precise irregularities. Furthermore, UTPA’s unified patch-level alignment elevates both detection and localization performance, demonstrating its complementary value.
State-Adjective Robustness Ablation
To verify our prompt’s robustness, we replace the original ‘good/damaged’ adjectives with several alternative pairs of similar meaning and retrain the model on MVTec-AD and VisA. As Table 4 shows, detection performance is virtually unchanged under every substitution, indicating high robustness to specific adjective choices. This resilience stems from our learnable prompt tokens, which autonomously capture normal and anomalous semantics, obviating the need for handcrafted text templates.
Figure 6: Top-k component ablation.
State words MVTec AD VisA
Pixel-level Image-level Pixel-level Image-level
Good/Damage (90.5, 87.1, 43.1) (91.2, 96.0, 92.0) (95.4, 92.2, 33.6) (88.5, 90.4, 84.8)
Normal/Abnormal (90.6, 87.5, 43.6) (91.1, 95.9, 92.3) (95.3, 91.9, 33.3) (88.3, 90.2, 84.5)
Perfect/Flawed (90.4, 87.2, 43.7) (91.7, 96.2, 92.5) (95.4, 92.0, 33.5) (88.3, 90.3, 84.5)
Flawless/Imperfect (90.4, 87.4, 43.3) (91.2, 95.9, 92.2) (95.4, 91.9, 33.1) (87.9, 90.2, 84.3)
Table 4: Ablation on different state words in prompts.
Top-k Ablation
To assess how the image-level representation is affected by the number of anomalous patches, we vary $K\in\{1,10,20,50,100,200\}$ and examine the AUROC performance in Figure 6. At $K=1$, the model is overly sensitive to local noise, yielding a low AUROC. As $K$ increases, performance steadily improves, demonstrating that aggregating multiple high-scoring anomalous patches enhances robustness. On MVTec AD, performance plateaus when $K$ is at 50 or 100; As for VisA, the optimum occurs at $K=50$. Beyond these points, additional patches introduce mostly normal regions, diluting the anomaly signals and reducing accuracy. Balanced across all metrics, we select the top 50 most anomalous patches as image-level descriptors.
6 Conclusion
We propose ViP2-CLIP, a universal ZSAD framework that detects anomalies in unseen categories without any target domain training samples. At its core, ViP-Prompt adaptively integrates global and local visual cues into text prompts, eliminating reliance on manual templates or class-name priors, achieving superior generalization and robustness in cross-modal alignment. In addition, the UTPA strategy optimizes a unified text–patch alignment across scales, sharpening both detection and localization performance. Extensive experiments on 15 mainstream datasets demonstrate that ViP2-CLIP consistently outperforms existing methods across various anomaly detection scenarios, especially when category labels are ambiguous or privacy-constrained, offering a more generalizable and scalable solution for ZSAD.
Limitations
UTPA is a simple yet powerful consistency alignment mechanism that optimizes a unified text–patch alignment for high-precision anomaly detection and localization. However, its current implementation only selects the better-performing Top-k from a small set of discrete candidates, while effective, it does not identify the dataset-specific optimal k. Since our primary goal is to demonstrate the effectiveness of optimizing global anomaly scores via aggregated anomalous patches within training-based CLIP models, further work on more advanced feature aggregation strategies is needed to explore.
Ethics Statement
Although ViP2-CLIP removes the reliance on explicit class labels, thereby minimizing exposure of sensitive category information, the anomaly maps can still expose private or proprietary details, particularly in medical imaging contexts. Practitioners must rigorously anonymize data, enforce strict access controls, and comply with all relevant privacy regulations to proprietary information.
Appendix A Motivation Statement
A.1 Robustness Analysis of Class-Name in Prompts
Methods like CLIP-AD (Chen et al., 2024), AdaCLIP (Cao et al., 2024), FILo (Gu et al., 2024), and KANOCLIP (Li et al., 2025) enhance alignment by embedding class names into prompts, but this introduces a critical vulnerability: CLIP’s segmentation quality is highly sensitive to label phrasing. Specifically, we conducted synonym-swap experiments on the training-free WinCLIP model to evaluate how CLIP’s inherent segmentation capability is influenced by the class-name variations. As shown in Figure 7, by replacing class names with semantically equivalent variants (e.g., ‘zipper’ to ‘zip’ or ‘fastener’), we observed a striking inconsistency: pixel-level segmentation metrics can vary by up to 10 percentage points, while image-level classification metrics remains stable, respectively. We attribute this instability to CLIP’s classification-oriented pre-training, which enforces robust global alignment but yields fragile local alignment.
To further evaluate whether training-based modality alignment can alleviate this sensitivity, we repeated the synonym-swap experiments on recent work, AdaCLIP. As shown in Figure 8, although AdaCLIP’s pixel-level F1 stays stable, its AUROC and PRO fluctuate with label swaps; in particular, changing ‘carpet’ to ‘floor mat’ triggers a 47% swing in PRO. This dramatic fluctuation shows that, even with fine-tuning on auxiliary data, the limited scale of anomaly-specific training cannot effectively mitigate CLIP’s segmentation sensitivity to class-name perturbations. These findings underscore the necessity of label-agnostic prompt designs for truly robust ZSAD.
Appendix B Experimental Details
B.1 Implementation Details
We use the publicly released CLIP (ViT-L/14@336px) model as the frozen feature extractor. In all experiments, the learnable prompt length is set to 10, of which 3 tokens fuse with the global visual feature. In UTPA, we set $Top\mbox{-}k=50$ and perform modal alignment at layers $\{6,12,18,24\}$. During training, input images are resized to $518\times 518$, we optimize with Adam at a learning rate of $1\times 10^{-3}$ for 10 epochs with a batch size of 8. All experiments are implemented in PyTorch 2.6.0 on a single NVIDIA L20 (42 GB).
Figure 7: WinCLIP’s performance of different Class Names in Prompts.
Figure 8: AdaCLIP’s performance of different Class Names in Prompts.
We conduct comprehensive evaluations across 15 publicly available benchmarks spanning industrial and medical domains. Specifically, we fine-tune on MVTec AD’s test set and perform zero-shot evaluation on all other datasets; for MVTec AD, we fine-tune on VisA’s test split. We adopt standard metrics for evaluation: Area Under the ROC Curve (AUC), Average Precision (AP), and F1-score (F1) for image-level anomaly detection; AUC, Per-Region Overlap (PRO), and F1 for pixel-level segmentation.
Domain Dataset Category Modalities $|\mathcal{C}|$ Normal and anomalous samples
Industrial MVTec AD Obj & texture Photography 15 (467, 1258)
VisA Obj Photography 12 (962, 1200)
MPDD Photography 6 (176, 282)
BTAD Photography 3 (451, 290)
KSDD Photography 1 (286, 54)
DAGM Texture Photography 10 (6996, 1054)
DTD-Synthetic Photography 12 (357, 947)
Medical ISIC Skin Photography 1 (0, 379)
CVC-ClinicDB Colon Endoscopy 1 (0, 612)
CVC-ColonDB Endoscopy 1 (0, 380)
Kvasir Endoscopy 1 (0, 1000)
Endo Endoscopy 1 (0, 200)
HeadCT Brain Radiology (CT) 1 (100, 100)
BrainMRI Radiology (MRI) 1 (98, 155)
Br35H Radiology (MRI) 1 (1500, 1500)
Table 5: Statistics of the datasets.
B.2 Baselines
To validate the superiority of ViP2-CLIP, we compare against the following state-of-the-art methods:
  • CLIP (Radford et al., 2021): We adapt CLIP for anomaly detection by using two prompt templates—‘a photo of a normal [cls]’ and ‘a photo of an abnormal [cls]’, where [cls] is the category class-name. We compute image-level anomaly scores via cosine similarity between global visual features and text embeddings. For pixel-level segmentation, we extend this scoring to the last-layer patch embeddings.
  • WinCLIP (Jeong et al., 2023): WinCLIP is the first CLIP-based ZSAD method, which uses multiple manually designed prompts and a multi-window feature extraction strategy. All parameters are kept the same as in their paper.
  • APRIL-GAN(Chen et al., 2023): APRIL-GAN extends WinCLIP by incorporating learnable linear projection layers to enrich local visual feature representations. All experiments are conducted using the official weights and adhere strictly to the original implementation protocols.
  • AnomalyCLIP (Zhou et al., 2024): AnomalyCLIP introduces object-agnostic learnable prompts and a DPAM mechanism for stronger local features’ modelling. All experiments are conducted using publicly released official model weights.
  • AdaCLIP (Cao et al., 2024): AdaCLIP incorporates hybrid prompts into both text and image encoders with an HSF module for improved cross-modal fusion. As AdaCLIP adopts an evaluation protocol distinct from ours, it trains on MVTec and CVC-ColonDB and performs zero-shot testing on the remaining datasets. To ensure fair and consistent comparison, we reimplement AdaCLIP under our unified evaluation protocol. All parameters are kept the same as in their paper.
Task Category Datasets PatchCore RD4AD ViP2-CLIP
Image-level (AUROC, AP) Obj &texture MVTec AD (99.0, 99.7) (98.7, 99.4) (91.2, 96.0)
Obj VisA (94.6, 95.9) (95.3, 95.7) (88.5, 90.4)
MPDD (94.1, 96.3) (91.6, 93.8) (79.7, 84.5)
BTAD (93.2, 98.6) (93.8, 96.8) (95.0, 98.4)
Texture DAGM (92.7, 81.3) (92.9, 79.1) (98.5, 94.3)
Pixel-level (AUROC, PRO) Obj &texture MVTec AD (98.1, 92.8) (97.8, 93.6) (90.5, 87.1)
Obj VisA (98.5, 92.2) (98.4, 91.2) (95.4, 92.2)
MPDD (98.8, 94.9) (98.4, 95.2) (97.2, 92.6)
BTAD (97.4, 74.4) (97.5, 75.1) (95.6, 86.1)
Texture DAGM (95.9, 87.9) (96.8, 91.9) (97.5, 95.2)
Table 6: Comparison of ZSAD performance between ViP2-CLIP and SOTA full-shot methods.Results of methods with * are copied from the AnomalyCLIP paper.The best performance is highlighted in red, and the second is highlighted in blue.
B.3 Datasets
We evaluate on 7 industrial datasets (MVTec AD, VisA, MPDD, BTAD, KSDD, DAGM, DTD-Synthetic) and 8 medical datasets (HeadCT, BrainMRI, Br35H, ISIC, CVC-ClinicDB, CVC-ColonDB, Endo, Kvasir), for a total of 15 benchmarks. All methods are trained and tested solely on each dataset’s test split. The dataset statistics are given in Table 5. We apply OpenCLIP’s default normalization to all images and resize them to $(518,518)$ to obtain suitably scaled feature maps.
Appendix C Additional Experiments and Analysis
C.1 Comparison with Full-shot Methods
In this section, we compare ViP2-CLIP against two state-of-the-art full-shot methods, PatchCore (Roth et al., 2022) and RD4AD (Deng and Li, 2022), on five industrial datasets for which training samples are available. Results are presented in Table 6. Despite not using any target domain training samples, ViP2-CLIP achieves detection and segmentation performance on par with or superior to these fully supervised approaches, with particularly strong gains on the BTAD and DAGM datasets. This demonstrates that ViP2-CLIP’s visual-perception prompt mechanism adaptively generates fine-grained normal and anomalous text descriptions that fully unleash CLIP’s ZSAD performance, breaking the dependency of traditional methods on in-domain training data and exhibits excellent generalization.
Model Training Time (h) Inference Time (ms)
AnomalyCLIP 1.02 90.72 $\pm$ 0.16
AdaCLIP 2.24 134.20 $\pm$ 0.39
ViP2-CLIP 0.54 48.69 $\pm$ 0.19
Table 7: Training and inference efficiency comparison. The best performance is highlighted in red.
C.2 Computational Efficiency Analysis
Table 7 compares training cost and per-image inference latency for the three strongest baselines. AnomalyCLIP’s DPAM rewrites some self-attention weights, stretching fine-tuning to 1.02 h and inference to 90.7 ms per image, while AdaCLIP’s extra learnable tokens in both encoders further burden the model. In contrast, ViP2-CLIP only attaches two lightweight adapters outside the frozen CLIP backbone, finishing training in 0.54 h and running inference in 48.7 ms per image, thus delivering the fastest runtime and lowest compute footprint among the three methods.
C.3 Targeted Fine-tuning on Medical Data
Although ViP2-CLIP achieves strong results on industrial benchmarks, its performance on medical images is relatively weaker, partly due to the domain gap between industrial auxiliary data and medical target domains. To investigate whether fine-tuning on medical auxiliary data can bridge this gap, we design a cross-domain fine-tuning protocol: we train on MVTec AD and CVC-ColonDB test split and test on the remaining medical datasets; for CVC-ColonDB, we use VisA and CVC-ClinicDB test sets as auxiliary data. We compare against AnomalyCLIP under the same training strategy. As shown in Table 8, introducing medical auxiliary data significantly boosts detection performance, particularly on CVC-ClinicDB and Endo datasets, highlighting the critical role of auxiliary data in cross-domain generalization. Overall, ViP2-CLIP outperforms AnomalyCLIP under medical fine-tuning, demonstrating superior cross-domain anomaly detection capability.
Task Datasets $|\mathcal{C}|$ AnomalyCLIP ViP2-CLIP
Image-level (AUROC, AP, F1) HeadCT 1 (94.3, 94.0, 90.5) (96.5, 96.5, 92.8)
BrainMRI 1 (94.7, 95.7, 92.5) (96.7, 97.6, 94.9)
Brain35H 1 (96.5, 96.7, 92.6) (97.1, 97.5, 93.4)
AVERAGE - (95.2, 95.5, 91.9) (96.8, 97.2, 93.7)
Pixel-level (AUROC, PRO, F1) ISIC 1 (87.8, 75.6, 69.7) (91.7, 82.9, 75.3)
CVC-ColonDB 1 (87.8, 76.9, 48.5) (87.7, 86.2, 53.5)
CVC-ClinicDB 1 (88.8, 76.1, 51.7) (92.2, 84.3, 60.8)
Endo 1 (89.0, 73.5, 58.1) (92.3, 82.8, 65.2)
Kvasir 1 (85.6, 50.0, 55.1) (90.5, 56.9, 63.4)
AVERAGE - (87.8, 70.4, 56.6) (90.9, 78.6, 63.6)
Table 8: ZSAD performance on medical images after fine-tuning by medical-domain datasets. The best performance is highlighted in red.
Figure 9: Prompt Length ablation.
C.4 Prompt Length Ablation
Next, based on VisA dataset, we vary the total prompt length to assess its effect on ZSAD performance, focusing on both the number of static learnable tokens $M$ and dynamic tokens $J$, which fuse with global visual context. Figure 9 shows that increasing length does not always improve performance: too many learnable tokens introduce redundancy and risk overfitting, harming precision and generalization capacity. To balance between semantic expressiveness and model robustness, we set the prompt length to 10 by default, comprising 3 dynamic tokens infused with global visual semantics and 7 static tokens modelling generic normal and anomalous patterns.
Appendix D Additional Quantitative and Qualitative Results
D.1 Fine-grained Subset Performance
Tables 920 report detailed ZSAD performance at the subset level on MVTec AD and VisA datasets, illustrating ViP2-CLIP’s ability to handle fine-grained variations within each category.
D.2 Visualization
We visualize pixel-wise anomaly maps in Figure 1023, demonstrating ViP2-CLIP’s strong segmentation capability and cross-domain generalization. Industrial examples include capsule, grid, leather, screw and hazelnut from MVTec AD; candle, cashew, macaroni, pcb and pipe fryum from VisA; brackets and metal plates from MPDD; and unknown products from BTAD. Medical examples cover melanoma detection in ISIC, and colorectal polyp detection in CVC-ColonDB and CVC-ClinicDB.
Object name CLIP WinCLIP APRIL-GAN AdaCLIP AnomalyCLIP ViP2-CLIP
Candle 32.7 87.0 97.8 98.9 98.8 99.0
Capsules 44.8 80.0 97.5 98.7 95.0 98.3
Cashew 21.8 84.8 85.8 92.1 93.8 91.4
Chewinggum 37.8 95.4 99.5 99.6 99.3 99.7
Fryum 26.0 87.7 91.9 94.5 94.6 92.6
Macaroni1 52.9 50.3 98.8 99.2 98.3 99.5
Macaroni2 70.4 44.7 97.8 98.4 97.6 98.7
Pcb1 61.9 38.6 92.8 93.7 94.0 90.1
Pcb2 27.4 58.7 89.8 90.8 92.4 92.0
Pcb3 71.7 76.0 88.2 88.5 88.4 92.2
Pcb4 49.2 91.4 94.5 96.1 95.7 96.4
Pipe_fryum 26.9 83.6 96.0 96.0 98.2 95.2
Mean 43.6 73.2 94.2 95.5 95.5 95.4
Table 9: Fine-grained performance comparison for Pixel-level AUROC on VisA.
Object name CLIP WinCLIP APRIL-GAN AnomalyCLIP AdaCLIP ViP2-CLIP
Candle 7.9 77.7 92.3 96.5 62.2 96.7
Capsules 12.0 39.4 86.1 78.9 38.3 93.5
Cashew 0.3 78.4 91.5 91.9 57.5 94.9
Chewinggum 9.1 69.6 87.5 90.9 55.8 96.8
Fryum 6.1 74.4 89.4 86.9 52.7 94.8
Macaroni1 25.9 24.8 93.0 89.8 65.6 96.4
Macaroni2 32.6 8.0 82.0 84.0 62.9 89.5
Pcb1 18.4 20.7 87.3 80.7 47.7 91.4
Pcb2 8.2 20.7 75.4 78.9 50.2 80.0
Pcb3 27.2 43.8 77.2 76.8 41.8 84.3
Pcb4 11.6 74.5 86.6 89.4 63.9 90.7
Pipe_fryum 9.1 80.3 90.9 96.2 83.2 97.2
Mean 14.0 51.0 86.6 86.7 56.8 92.2
Table 10: Fine-grained performance comparison for Pixel-level AUPRO on VisA.
Object name CLIP WinCLIP APRIL-GAN AnomalyCLIP AdaCLIP ViP2-CLIP
Candle 0.3 8.9 39.4 37.8 42.1 42.4
Capsules 1.1 4.2 49.1 37.8 54.1 50.7
Cashew 2.2 9.6 22.7 25.8 35.1 23.8
Chewinggum 1.1 31.6 78.5 61.0 78.2 74.4
Fryum 4.5 16.2 29.5 30.3 32.7 25.6
Macaroni1 0.1 0.1 35.3 23.7 32.7 38.9
Macaroni2 0.1 0.1 13.9 5.1 14.1 9.7
Pcb1 2.3 0.9 12.2 12.7 13.7 13.4
Pcb2 0.4 1.5 23.3 15.8 32.5 23.4
Pcb3 1.2 2.1 21.9 9.3 35.8 25.7
Pcb4 1.5 24.6 31.0 34.7 43.0 39.8
Pipe_fryum 2.5 8.3 30.4 45.5 29.5 35.9
Mean 1.5 9.0 32.3 28.3 37.0 33.6
Table 11: Fine-grained performance comparison for Pixel-level F1 on VisA.
Object name CLIP WinCLIP APRIL-GAN AnomalyCLIP AdaCLIP ViP2-CLIP
Candle 55.2 94.8 82.6 80.9 96.0 90.4
Capsules 61.9 79.4 62.3 82.8 87.0 93.8
Cashew 68.3 91.2 86.6 76.0 88.1 96.4
Chewinggum 59.4 95.5 96.4 97.2 93.9 97.9
Fryum 48.4 73.6 93.8 92.7 91.8 88.5
Macaroni1 61.5 79.0 69.3 86.7 84.2 86.3
Macaroni2 52.3 67.1 65.7 72.2 67.5 70.8
Pcb1 66.1 72.1 51.0 85.2 89.2 90.8
Pcb2 64.4 47.0 71.4 62.0 84.3 78.8
Pcb3 50.7 63.9 66.9 61.7 78.9 78.3
Pcb4 64.8 74.2 94.7 93.9 95.3 97.5
Pipe_fryum 68.9 67.9 89.2 92.3 90.3 92.4
Mean 60.2 75.5 77.5 82.0 87.2 88.5
Table 12: Fine-grained performance comparison for Image-level AUROC on VisA.
Object name CLIP WinCLIP APRIL-GAN AnomalyCLIP AdaCLIP ViP2-CLIP
Candle 56.1 95.4 86.0 82.6 96.6 92.6
Capsules 74.9 87.9 74.5 89.4 92.5 96.7
Cashew 80.9 96.0 93.8 89.3 94.9 98.4
Chewinggum 77.3 98.2 98.4 98.8 97.5 99.1
Fryum 70.4 86.9 97.1 96.6 96.3 94.5
Macaroni1 60.3 80.0 67.4 85.5 83.1 87.9
Macaroni2 49.9 65.1 64.8 70.8 67.7 71.1
Pcb1 69.9 73.0 55.3 86.7 89.4 91.0
Pcb2 62.5 46.1 73.4 64.4 85.6 78.9
Pcb3 50.6 63.1 70.4 69.4 82.4 81.6
Pcb4 59.6 70.1 94.8 94.3 95.5 96.8
Pipe_fryum 82.1 82.1 94.5 96.3 95.1 96.3
Mean 66.2 78.7 80.9 85.3 89.7 90.4
Table 13: Fine-grained performance comparison for Image-level AP on VisA.
Object name CLIP WinCLIP APRIL-GAN AnomalyCLIP AdaCLIP ViP2-CLIP
Candle 67.4 90.6 77.9 75.6 89.5 83.0
Capsules 76.9 80.5 78.0 82.2 85.0 90.8
Cashew 80.2 88.9 85.4 80.3 86.8 94.0
Chewinggum 80.0 93.8 93.2 94.8 91.5 96.4
Fryum 80.0 80.0 91.5 90.1 88.2 88.2
Macaroni1 71.5 74.2 70.8 80.4 80.4 79.6
Macaroni2 66.7 68.8 69.3 71.2 68.8 71.3
Pcb1 68.1 70.2 66.9 78.8 81.5 85.0
Pcb2 68.4 67.1 69.1 67.8 77.5 74.1
Pcb3 66.4 67.6 66.7 66.4 73.6 73.3
Pcb4 69.6 75.7 87.3 87.8 89.7 92.4
Pipe_fryum 80.8 80.3 88.1 89.8 89.0 89.5
Mean 73.0 78.2 78.7 80.4 83.5 84.8
Table 14: Fine-grained performance comparison for Image-level F1 on VisA.
Object name CLIP WinCLIP VAND AnomalyCLIP AdaCLIP ViP2-CLIP
Carpet 18.0 90.9 98.4 98.8 98.6 99.2
Bottle 19.6 85.7 83.5 90.4 92.8 89.4
Hazelnut 27.6 95.7 96.1 97.2 98.6 96.7
Leather 12.7 95.5 99.1 98.6 99.3 99.2
Cable 44.1 61.3 72.3 78.9 76.6 73.2
Capsule 58.0 87.0 92.0 95.8 94.3 94.4
Grid 11.8 79.4 95.8 97.3 91.1 97.7
Pill 45.6 72.7 76.2 91.8 86.7 87.5
Transistor 42.2 83.7 62.4 70.8 63.5 64.6
Metal Nut 33.3 49.3 65.5 74.6 68.2 76.8
Screw 72.3 91.1 97.8 97.5 97.8 98.7
Toothbrush 26.0 86.2 95.8 91.9 97.2 93.4
Zipper 52.5 91.7 91.1 91.3 95.7 97.0
Tile 34.6 79.1 92.7 94.7 89.5 93.7
Wood 36.1 85.1 95.8 96.4 94.0 95.7
Mean 35.6 82.3 87.6 91.1 89.6 90.5
Table 15: Fine-grained performance comparison for Pixel-level AUROC on MVTec.
Object name CLIP WinCLIP VAND AnomalyCLIP AdaCLIP ViP2-CLIP
Carpet 6.2 66.3 48.5 90.0 38.1 97.9
Bottle 0.3 69.9 45.6 80.8 39.0 83.5
Hazelnut 4.7 81.4 70.3 92.5 19.4 93.1
Leather 1.4 86.0 72.4 92.2 57.0 98.8
Cable 8.8 39.4 25.7 64.0 43.1 66.5
Capsule 31.6 63.7 51.3 87.6 60.3 92.7
Grid 0.2 49.3 31.6 75.4 57.9 89.5
Pill 5.3 66.9 65.4 88.1 40.5 93.1
Transistor 8.6 45.5 21.3 58.2 27.1 55.1
Metal Nut 0.9 39.6 38.4 71.1 63.9 79.8
Screw 48.5 70.2 67.1 88.0 16.1 94.2
Toothbrush 2.3 67.9 54.5 88.5 58.9 88.8
Zipper 20.8 72.1 10.7 65.4 18.1 89.1
Tile 6.3 54.5 26.7 87.4 25.7 89.2
Wood 12.9 56.3 31.1 91.5 2.4 95.5
Mean 10.6 61.9 44.0 81.4 37.8 87.1
Table 16: Fine-grained performance comparison for Pixel-level AUPRO on MVTec.
Object name CLIP WinCLIP VAND AnomalyCLIP AdaCLIP ViP2-CLIP
Carpet 3.2 33.9 65.7 57.0 59.5 65.5
Bottle 10.9 49.4 53.4 51.6 32.2 52.3
Hazelnut 4.2 39.1 50.5 47.6 34.2 49.9
Leather 1.3 30.8 50.0 33.2 66.6 43.0
Cable 5.6 12.2 23.9 18.9 37.0 24.1
Capsule 4.8 14.3 33.1 31.0 63.3 34.4
Grid 1.4 13.7 40.8 32.0 49.8 41.3
Pill 7.4 11.8 27.7 35.5 31.2 26.0
Transistor 9.2 27.0 19.0 18.8 31.7 16.0
Metal Nut 21.0 23.8 28.0 33.1 28.6 35.1
Screw 6.4 11.3 41.7 33.4 67.4 46.4
Toothbrush 3.0 10.5 48.1 29.0 47.8 31.9
Zipper 5.2 27.8 40.5 45.0 18.0 53.2
Tile 13.2 30.8 66.5 64.9 52.9 68.0
Wood 7.4 35.4 60.3 55.2 55.6 58.7
Mean 6.9 24.8 43.3 39.1 45.1 43.1
Table 17: Fine-grained performance comparison for Pixel-level F1 on MVTec.
Object name CLIP WinCLIP VAND AnomalyCLIP AdaCLIP ViP2-CLIP
Carpet 86.9 99.3 99.4 100 99.9 99.9
Bottle 20.5 98.6 91.9 88.7 97.9 93.7
Hazelnut 53 92.3 89.4 97.2 97.2 93.5
Leather 98.5 100 99.7 99.8 99.9 100
Cable 51.4 85 88.1 70.3 65.5 75.3
Capsule 60.7 68.6 79.9 89.5 86.5 90.2
Grid 77 99.2 86.5 97.8 99.4 99.7
Pill 61.2 81.5 80.9 81.1 84.6 84.6
Transistor 45.6 89.1 80.8 93.9 82.2 79.9
Metal Nut 68.2 96.2 68.4 92.4 75.9 77.3
Screw 62.2 71.6 85.1 82.1 97.8 90
Toothbrush 23.6 85.3 53.6 85.3 93.3 94.2
Zipper 90 91.2 89.6 98.4 96.4 92.9
Tile 98.9 99.9 99.9 100 99.9 98.5
Wood 99.6 97.6 99 96.9 99.0 98.6
Mean 66.5 90.4 86.1 91.6 90.1 91.2
Table 18: Fine-grained performance comparison for Image-level AUROC on MVTec.
Object name CLIP WinCLIP VAND AnomalyCLIP AdaCLIP ViP2-CLIP
Carpet 95.6 97.8 99.8 100 99.9 100
Bottle 65 97.6 97.6 96.8 99.4 98.1
Hazelnut 67.2 88.6 94.6 98.5 98.4 96.8
Leather 99.5 100 99.9 99.9 99.9 100
Cable 66.2 89.8 92.9 81.7 81.8 84.5
Capsule 88.5 90.5 95.4 97.8 97.1 97.9
Grid 90.6 98.2 94.9 99.3 99.8 99.9
Pill 89.9 96.4 96.1 95.3 96.6 96.6
Transistor 46.3 84.9 77.5 92.1 82.7 80
Metal Nut 91.1 99.1 91.8 98.2 94.5 94.7
Screw 81.6 87.7 93.6 92.9 96.5 96.5
Toothbrush 60.7 94.5 71.6 93.9 97.7 97.5
Zipper 96.6 97.5 97.1 99.5 99.1 98.2
Tile 99.5 100 100 100 100 99.5
Wood 99.9 99.3 99.7 99.2 99.7 99.6
Mean 82.6 95.6 93.5 96.4 95.6 96
Table 19: Fine-grained performance comparison for Image-level AP on MVTec.
Object name CLIP WinCLIP VAND AnomalyCLIP AdaCLIP ViP2-CLIP
Carpet 89.4 97.8 98.3 99.4 99.4 99.4
Bottle 86.3 97.6 92.1 90.9 92.4 92.4
Hazelnut 77.8 88.6 87 92.6 91.2 91.2
Leather 96.7 100 98.9 99.5 99.5 100
Cable 76 84.8 84.5 77.4 76.0 80.2
Capsule 90.5 93.5 91.5 91.7 93.1 93
Grid 88.4 98.2 89.1 97.3 98.3 99.1
Pill 91.6 91.6 91.6 92.1 91.8 91.8
Transistor 57.1 80 73.9 83.7 75.9 71.8
Metal Nut 89.4 95.3 89.4 93.7 89.4 90.3
Screw 85.6 85.9 89.3 88.3 91.3 91.3
Toothbrush 83.3 88.9 83.3 90 93.5 93.5
Zipper 91.5 93.4 90.8 97.9 95.1 92.4
Tile 97 99.4 99.4 100 99.4 96.4
Wood 99.2 95.2 97.4 96.6 96.7 96.8
Mean 86.7 92.7 90.4 92.7 92.3 92
Table 20: Fine-grained performance comparison for Image-level F1 on MVTec.
Figure 10: Anomaly score maps for the data subset, capsule, in MVTec AD. The first row represents the input, the second row presents the segmentation results from ViP2-CLIP.
Figure 11: Anomaly score maps for the data subset, leather, in MVTec AD. The first row represents the input, the second row presents the segmentation results from ViP2-CLIP.
Figure 12: Anomaly score maps for the data subset, wood, in MVTec AD. The first row represents the input, the second row presents the segmentation results from ViP2-CLIP.
Figure 13: Anomaly score maps for the data subset, screw, in MVTec AD. The first row represents the input, the second row presents the segmentation results from ViP2-CLIP.
Figure 14: Anomaly score maps for the data subset, hazelnut, in MVTec AD. The first row represents the input, the second row presents the segmentation results from ViP2-CLIP.
Figure 15: Anomaly score maps for the data subset, candle, in VisA. The first row represents the input, the second row presents the segmentation results from ViP2-CLIP.
Figure 16: Anomaly score maps for the data subset, cashew, in VisA. The first row represents the input, the second row presents the segmentation results from ViP2-CLIP.
Figure 17: Anomaly score maps for the data subset, pcb, in VisA. The first row represents the input, the second row presents the segmentation results from ViP2-CLIP.
Figure 18: Anomaly score maps for the data subset, pipe fryum, in VisA. The first row represents the input, the second row presents the segmentation results from ViP2-CLIP.
Figure 19: Anomaly score maps for the data subset in MPDD. The first row represents the input, the second row presents the segmentation results from ViP2-CLIP.
Figure 20: Anomaly score maps for the data subset in BTAD. The first row represents the input, the second row presents the segmentation results from ViP2-CLIP.
Figure 21: Anomaly score maps for the data subset skin. The first row represents the input, the second row presents the segmentation results from ViP2-CLIP.
Figure 22: Anomaly score maps for the data subset colon. The first row represents the input, the second row presents the segmentation results from ViP2-CLIP.
Figure 23: Anomaly score maps for the data subset colon. The first row represents the input, the second row presents the segmentation results from ViP2-CLIP.